Find the quotient by 0.9 / 0.45
i. changing these numbers into fraction form,
applying the fraction procedure, and changing
the product back to decimal form. (I get 9/10 / 9/20 but I am not sure if that is correct and if so what to do next?)
ii. applying the decimal procedure to the numbers in
decimal form.
b. How do the two quotients compare? How do the
two procedures compare?

Answer by richwmiller(17219)

.9 9/10 also equals .90 90/100
.90/.45=2
the easiest way to do this is multiply both top and bottom by 100
90/45=2
90/100/45/100=90/100*100/45=90/45=2
your 9/10/9/20=9/10*20/9=20/10=2
but where did you get 9/20 and why?